CPI sells computer peripherals. At December 31, year 1, CPI’s inventory amounted to $620,000. During the first week in January, year 2, the company made only one purchase and one sale. These transactions were as follows. Jan. 2 Purchased 20 modems and 80 printers from Sharp. The total cost of these machines was $37,000, terms 3/10, n/60. Jan. 6 Sold 30 different types of products on account to Pace Corporation. The total sales price was $22,000, terms 5/10, n/90. The total cost of these 30 units to CPI was $12,540 (net of the purchase discount). CPI has a full-time accountant and a computer-based accounting system. It records sales at the gross sales price and purchases at net cost and maintains subsidiary ledgers for accounts receivable, inventory, and accounts payable. Required: b. Prepare journal entries to record these transactions, assuming that CPI uses a perpetual inventory system. c. Compute the balance in the Inventory account at the close of business on January 6. d. Prepare journal entries to record the two transactions, assuming that CPI uses a periodic inventory system. e. Compute the cost of goods sold for the first week of January assuming use of the periodic system. (Use your answer to part c as the ending inventory.) g. Compute the gross profit margin on the January 6 sales transaction.
Answer #1

b).

Date Account title Debit credit
Jan 2 Merchandise inventory 37000
Accounts payable 37000
Jan 6 Accounts receivable 22000
sales revenue 22000
To record sales]
cost of goods sold 12540
Merchandise inventory 12540

c)purchase ,net of discount = Gross purchase (1-discount)

                       = 37000(1-.03)

                       = 37000*.97

                       = 35890

Ending inventory =Beginning inventory +Net purchase-cost of sales

                   = 620000+35890-12540

                    = 643350

d)

Date Account title Debit credit
Jan 2 Purchase 37000
Accounts payable 37000
Jan 6 Accounts receivable 22000
sales revenue 22000
To record sales]

e)

Cost of goods sold =Beginning inventory + Net purchase -ending inventory

                         = 620000+35890-643350

                         = 12540

g)Gross margin =sales -cost of goods sold

                     = 22000-12540

                        = 9460